OpenStudy (anonymous):

There were 512 people at the talent show. Admission was $1.25 for adults and $ 0.50 for children. The total amount collected was $406.75. How many children attended the show

OpenStudy (anonymous):

make a parabola equation

OpenStudy (anonymous):

512= x+y 406.75=.5x+1.25y children= 311 adults 201